OpenStudy (anonymous):

Find the slope of the line between (-3, 4) and (-6, -8)

OpenStudy (lgbasallote):

\[\text{slope} = \frac{y_2 - y_1}{x_2 - x_1}\] \[\text{slope} = \frac{-8 - 4}{-6 - (-3)}\] do you have an answer?

OpenStudy (anonymous):

solpe=change in y/change in x =( 4-(-8))/(-3-(-6)=12/3=4